# Taxonomy and Trends in Reinforcement Learning for Robotics and Control Systems: A Structured Review

### v1

@@ I Introduction
Reinforcement Learning (RL), a subfield of machine learning, has emerged as a powerful paradigm for solving sequential decision-making problems [1, 2]. In RL, agents learn optimal behaviors through trial-and-error interactions with dynamic environments, guided by feedback in the form of rewards or penalties [3]. This learning strategy is particularly well-suited for robotics and control systems, where agents must operate under uncertainty, interact with physical environments, and adapt to evolving tasks and conditions [4, 5].
Traditional control methods, such as Proportional-Integral-Derivative (PID) controllers, Linear Quadratic Regulators (LQR), and Model Predictive Control (MPC), have long been the cornerstone of robotics [6, 7]. These methods rely on accurate system models and often require manual tuning to perform well in complex, nonlinear, or changing environments [8]. Reinforcement learning, by contrast, offers the potential for model-free or partially model-based control strategies that can automatically adapt policies through direct interaction, even when the system dynamics are unknown or too complex to model precisely [9].
@@ III Taxonomy of Reinforcement Learning Applications in Robotics
The first component of the taxonomy categorizes RL applications based on robot capabilities and interaction dynamics [38]. In the domain of single-robot operation, key competencies are grouped into locomotion, navigation, stationary manipulation, and mobile manipulation. Locomotion refers to the robot’s fundamental ability to move, typically through legs, wheels, or tracks [37]. Navigation builds upon locomotion by introducing goal-directed movement in structured or unstructured environments [39, 40, 41]. In parallel, manipulation tasks may occur in stationary settings (e.g., robotic arms in manufacturing) or in mobile contexts, where navigation and manipulation are integrated [42].
@@ IV-D Industrial Automation
  • Adaptive Quality Control: Inspection robots trained with RL can adapt their scanning behavior based on defect likelihood particularly with virtual reality [57, 58, 59].

### v2

@@ I Introduction
Reinforcement Learning (RL), a subfield of machine learning, has emerged as a powerful paradigm for solving sequential decision-making problems [1, 2]. In RL, agents learn optimal behaviors through trial-and-error interactions with dynamic environments, guided by feedback in the form of rewards or penalties [3]. This learning strategy is particularly well-suited for robotics and control systems, where agents must operate under uncertainty, interact with physical environments, and adapt to evolving tasks and conditions [4].
Traditional control methods, such as Proportional-Integral-Derivative (PID) controllers, Linear Quadratic Regulators (LQR), and Model Predictive Control (MPC), have long been the cornerstone of robotics [6]. These methods rely on accurate system models and often require manual tuning to perform well in complex, nonlinear, or changing environments [8]. Reinforcement learning, by contrast, offers the potential for model-free or partially model-based control strategies that can automatically adapt policies through direct interaction, even when the system dynamics are unknown or too complex to model precisely [9].
@@ III Taxonomy of Reinforcement Learning Applications in Robotics
The first component of the taxonomy categorizes RL applications based on robot capabilities and interaction dynamics [38]. In the domain of single-robot operation, key competencies are grouped into locomotion, navigation, stationary manipulation, and mobile manipulation. Locomotion refers to the robot’s fundamental ability to move, typically through legs, wheels, or tracks [37]. Navigation builds upon locomotion by introducing goal-directed movement in structured or unstructured environments [41]. In parallel, manipulation tasks may occur in stationary settings (e.g., robotic arms in manufacturing) or in mobile contexts, where navigation and manipulation are integrated [42].
@@ IV-D Industrial Automation
  • Adaptive Quality Control: Inspection robots trained with RL can adapt their scanning behavior based on defect likelihood particularly with virtual reality [57].